# Fast Perfekt: Regression-based refinement of fast simulation

### v1

@@ Abstract
The availability of precise and accurate simulation is a limiting factor for interpreting and forecasting data in many fields of science and engineering. Often, one or more distinct simulation software applications are developed, each with a relative advantage in accuracy or speed. The quality of insights extracted from the data stands to increase if the accuracy of faster, more economical simulation could be improved to parity or near parity with more resource-intensive but accurate simulation. We present Fast Perfekt, a machine-learned regression that employs residual neural networks to refine the output of fast simulations. A deterministic morphing model is trained using a unique schedule that makes use of the ensemble loss function MMD, with the option of an additional pair-based loss function such as the MSE. We explore this methodology in the context of an abstract analytical model and in terms of a realistic particle physics application featuring jet properties in hadron collisions at the CERN Large Hadron Collider. The refinement makes maximum use of domain knowledge, and introduces minimal computational overhead to production.
@@ 1 Introduction
Significant computing and scientific benefits could be realized if fastsim were made highly accurate while maintaining its speed advantage. Recent approaches have been developed in particle physics to render data emulated by fastsim more accurate. For example, the DCTR [9] approach weights simulated data points such that their distributions more closely agree with target (fullsim) data, correcting distributions of features and their correlations. Limitations can arise, however, because the weights are specific to the underlying process, the support is limited to the domain of the input, and the weights reduce the statistical precision of the fastsim data. Other methods employ a Wasserstein metric or integral loss function to update simulated features via a mapping  [10, 11, 12], the first using generative methods and the last two being deterministic mappings. These approaches can be effective in refining simulation without introducing weights. However, these methods typically do not exploit all relevant information present in the training samples, in particular the object-to-object correlations between fastsim and fullsim. Moreover, while probability density functions (PDFs) of the refined feature(s) may agree well with the target, the transformation may not be unique and can lead to degraded correlations with features not directly used by the network.
@@
The concept and procedure for training the network are described in Section 2. The methods are demonstrated in Section 3 using an example based on an abstract analytical data set. Fast Perfekt is then applied in a realistic physics application in Section 4 featuring jet substructure variables in LHC experiments. Section 5 summarizes the main findings, impact and limitations.
@@ 2 Method
Let $\mathbf{x^{\prime}}$ be a set of fastsim features to be refined. These features, or correlations among them, presumably exhibit inaccuracies with respect to the fullsim features $\mathbf{x}$. Each element of $\mathbf{x^{\prime}}$ is taken as input to the refiner network along with any conditioning variables, such as the GT values $\mathbf{g}$ of the given object, together defining the vector $\intercal\mathbf{a^{\prime}}=(\mathbf{g},\mathbf{x^{\prime}})^{\intercal}$. Additional features $\mathbf{h^{\prime}}$ are hidden, either unavailable or not chosen among the set to be refined, but may also be important for downstream or upstream analysis. The network outputs refined features $\mathbf{\hat{x}}$ whose properties and correlations to the hidden features $\mathrm{corr}(\mathbf{\hat{x}},\mathbf{h^{\prime}})$ are estimators of the fullsim features $\mathrm{x}$ and correlations $\textbf{corr}(\mathbf{x},\mathbf{h})$, respectively. We note that the network goal is to render fastsim features more fullsim-like, and not more like the GT values. A simple schematic of the sampling and training procedure is given in Figure 1.
@@ 2.2 Loss functions
The biased and unbiased MMD are highly correlated, but $\text{MMD}_{\mathrm{u}}$ has properties that better facilitate convergence in the first training stage, due to its having a well-defined expectation value of 0 for two i.i.d. data sets, which can easily be fulfilled as a constraint in the MDMM. Then, $\text{MMD}_{\mathrm{b}}$ is better suited for refining the finest details because of its well-defined lower boundary at 0 corresponding to two ensembles in perfect agreement. Minimizing the MSE alone would result in significant biases due to regression to the mean. However, because it encodes information about correlations among visible and hidden features, minimizing the MSE conditionally while constraining $\text{MMD}_{\mathrm{u}}$ to 0 serves to protect or restore correlations.
@@ 2.3 Training
The constrained minimization is implemented using the modified differential method of multipliers (MDMM) [17], which identifies saddle points in the space of the network parameters and Lagrange multiplier $\lambda$. The Lagrangian is
[…]
@@ 3.1 Data set
Taking $m=250000$ to be the total number of data samples $\mathbf{g}=(x^{\mathrm{GT}},h^{\mathrm{GT}})^{T}$ used, a large population in the GT distribution is sampled from $G_{L}$ with $0.85N$ samples $\mathbf{g}_{0},\dots,\mathbf{g}_{0.85m-1}$ and a small population is sampled from $G_{s}$ with $0.15m$ samples $\mathbf{g}_{0.85m},\dots,\mathbf{g}_{m}$, without loss of generality.
@@ 4 Application to collider physics
The refinement protocol using the 2-stage training scheme introduced above is applied to a realistic example based on high-energy physics data collected at the LHC. The data set is described, choices regarding the training process are discussed, and the results of the refinement application follow.
@@ 4.1 Data set
Jets are notoriously difficult to model because their production mechanisms are only partly calculable with perturbative methods and because the energetic response of the detectors is a complicated function of the detector geometry, event activity, and radiation exposure. Considerable effort has been made to accurately model jets in state-of-the-art fullsim models such as Geant4-based [2] applications of ATLAS and CMS. Fullsim processing time per event ranges from 10s of second to a few minute per event, depending on the underlying physics process. In fastsim programs, such as ATLAS AtlFast3 [3], CMS FastSim [4, 5], the processing time is only a few seconds, but due to simplified assumptions in the calorimetry and tracking detectors, mismodeling of jets is often compounded. In parametric simulators, such as the Delphes framework [6], the processing time per event ranges in the milliseconds, but jet mismodeling arises due to simplified energy response parametrizations.
@@
We consider a set of jet substructure observables called N-subjettiness $\tau_{N}$ [23]. Ratios such as $\tau_{2}/\tau_{1}$ provide information on the compositeness of the jets to infer the unobservable particles from which they originated. A GT data set is generated based on proton-proton collisions at $\sqrt{s}=13\,\mathrm{TeV}$ using the Pythia 8.1 software package [24], consisting of events with a pair of top quarks in the final state. These events are then processed twice in parallel using Delphes, once with the default CMS detector implementation resulting in a fullsim data set, and once with a “flawed" implementation yielding the data set we treat as the fastsim. The flaw is introduced by setting the $\beta$ parameter [23] to 0.9 rather than 1.0, inducing an angle-dependent bias in the N-subjettiness, and setting the energy resolution of the electromagnetic calorimeter to a constant of 1% rather than the default values, which depend on transverse momentum $p_{T}$ and pseudorapidity $\eta$ [25, 26].
@@ 4.2 Training
$x^{\prime}$ $\intercal\displaystyle=\begin{pmatrix}\frac{\tau_{2}}{\tau_{1}},&\frac{\tau_{3}}{\tau_{2}},&\frac{\tau_{4}}{\tau_{3}}\end{pmatrix}^{\intercal}$ (8)
$g$ $\intercal\displaystyle=\begin{pmatrix}\frac{\tau_{2}^{GT}}{\tau_{1}^{GT}},&\frac{\tau_{3}^{GT}}{\tau_{2}^{GT}},&\frac{\tau_{4}^{GT}}{\tau_{3}^{GT}}\end{pmatrix}^{\intercal}\,.$ (9)
@@ 5 Conclusions
We have introduced a regression-based procedure for refining the final analysis variables produced by a fast simulation application (fastsim) to render them in milliseconds highly accurate with respect to cost-intensive full simulation (fullsim). The refinement model, which is based on a residual neural network, is trained by minimizing an ensemble loss (MMD), which evaluates the similarity between two unbinned multi-dimensional distributions. As an additional option, a secondary pair-based loss (MSE) is used in a 2-stage training prescription: one stage that combines the MMD and MSE losses, and a second stage which uses only the MMD.
Two example problems were explored for the refinement application, each considering the accuracy of features that are available to the network as well as hidden features which are not. In the first example, a proxy fastsim and fullsim data set is generated based on a simplified analytical model, with the two samples synchronized at the level of the ground truth. We find that the single-stage MMD-only procedure performs well when evaluating the accuracy of the available feature; however, it falls short in describing the multidimensional target space which includes the hidden dimension. When introducing the 2-stage prescription, the accuracy of the refinement improves, and moreover the multidimensional domain is well described. A second example based on a realistic model of collider physics at the CERN LHC examines the performance of the two prescriptions with a larger number of available and hidden features. Here, the single-stage and 2-stage prescriptions lead to a similarly significant improvement in the accuracy of the fastsim.
Fast Perfekt makes comprehensive use of domain knowledge present in existing fastsim programs, minimizing the need for the network to learn most of the salient features and correlations present in the target data. The utility of the MMD as a loss function for sample regression has been demonstrated, and the MMD is deemed suitable for refinement problems that target real-world rather than simulated data. However, as with all integral loss functions, a downside of the MMD is the large multiplicity of unique network configurations corresponding to the global minimum. The two-stage Fast Perfekt prescription constrains the network to a highly unique configuration by conditionally minimizing the pair-based MSE loss, which increases the stability and overall accuracy of the refined simulation, rendering the two stage method ideal for tuning fastsim to match fullsim.

### v2

@@ Abstract
The availability of precise and accurate simulation is a limiting factor for interpreting and forecasting data in many fields of science and engineering. Often, one or more distinct simulation software applications are developed, each with a relative advantage in accuracy or speed. The quality of insights extracted from the data stands to increase if the accuracy of faster, more economical simulation could be improved to parity or near parity with more resource-intensive but accurate simulation. We present Fast Perfekt, a machine-learned regression to refine the output of fast simulations that employs residual neural networks. A deterministic morphing model is trained using a unique schedule that makes use of the ensemble loss function MMD, with the option of an additional pair-based loss function such as the MSE. We explore this methodology in the context of an abstract analytical model and in terms of a realistic particle physics application featuring jet properties in hadron collisions at the CERN Large Hadron Collider. The refinement makes maximum use of existing domain knowledge, and introduces minimal computational overhead to production.
@@ 1 Introduction
Significant computing and scientific benefits could be realized if fastsim were made highly accurate while maintaining its speed advantage. Recent approaches have been developed in particle physics to render data emulated by fastsim more accurate. For example, the DCTR [9] approach weights simulated data points such that their distributions more closely agree with target (fullsim) data, correcting distributions of features and their correlations. Limitations can arise, however, because the weights are specific to the underlying process, the support is limited to the domain of the input, and the weights reduce the statistical precision of the fastsim data. Other methods employ a Wasserstein metric or integral loss function to update simulated features via a mapping  [10, 11, 12], the first using generative methods and the last two being deterministic mappings. These approaches can be effective in refining simulation without introducing weights. However, these methods typically do not exploit all relevant information present in the training samples, in particular the object-to-object correlations between fastsim and fullsim. Moreover, while probability density functions (PDFs) of the refined feature(s) may agree well with the target, the transformation may not be unique and can lead to degraded correlations with features not directly used by the network. More generally, ML has been used to replace components of fast simulation like the modeling of calorimeter showers, such as the seminal work [13], with a broad review given in [14]; similar techniques have been applied in ATLAS fastsim [3]. There are also efforts to replace the entire simulation frameworks with generative normalizing flows in CMS [15].
@@
Initial concepts of this method were previously reported as an application to CMS fast simulation (FastSim) [17]. The concept and procedure for training the network are described in Section 2. The methods are demonstrated in Section 3 using an example based on an abstract analytical data set. Fast Perfekt is then applied in a realistic physics application in Section 4 featuring jet substructure variables in LHC experiments. Section 5 summarizes the main findings, impact and limitations.
@@ 2 Method
Let $\mathbf{x^{\prime}}$ be a set of fastsim features to be refined. These features, or correlations among them, presumably exhibit inaccuracies with respect to the fullsim features $\mathbf{x}$. Each element of $\mathbf{x^{\prime}}$ is taken as input to the refiner network along with any conditioning variables, such as the GT values $\mathbf{g}$ of the given object, together defining the vector $\intercal\mathbf{a^{\prime}}=(\mathbf{g},\mathbf{x^{\prime}})^{\intercal}$. Additional features $\mathbf{h^{\prime}}$ are hidden, either unavailable or not chosen among the set to be refined, but may also be important for downstream or upstream analysis. The network outputs refined features $\mathbf{\hat{x}}$ whose properties and correlations to the hidden features $\mathrm{corr}(\mathbf{\hat{x}},\mathbf{h^{\prime}})$ are estimators of the fullsim features $\mathbf{x}$ and correlations $\textbf{corr}(\mathbf{x},\mathbf{h})$, respectively. We note that the network goal is to render fastsim features more fullsim-like, and not more like the GT values; we also note that the network cannot refine the hidden variables themselves, but only correlations to them. A simple schematic of the sampling and training procedure is given in Figure 1.
@@ 2.2 Loss functions
The biased and unbiased MMD are highly correlated, but $\text{MMD}_{\mathrm{u}}$ has properties that better facilitate convergence in the first training stage, due to its having a well-defined expectation value of 0 for two i.i.d. data sets. This condition is reached via a constrained optimization, where we employ the modified differential method of multipliers (MDMM) [21], which identifies saddle points in the space of the network parameters and Lagrange multiplier $\lambda$. Then, $\text{MMD}_{\mathrm{b}}$ is better suited for refining the finest details because of its well-defined lower boundary at 0 corresponding to two ensembles in perfect agreement. Minimizing the MSE alone would result in significant biases due to regression to the mean. However, because it encodes information about correlations among visible and hidden features, minimizing the MSE conditionally while constraining $\text{MMD}_{\mathrm{u}}$ to 0 serves to protect or restore correlations.
@@ 2.3 Training
The Lagrangian used in the MDMM is
$\mathcal{L}(\mathbf{\theta},\lambda)$ $=\mathrm{MSE}(\mathbf{\theta})-\lambda\big{(}\varepsilon-\mathrm{MMD}_{\mathrm{u}}(\mathbf{\theta})\big{)}-\frac{\delta}{2}\big{(}\varepsilon-\mathrm{MMD}_{\mathrm{u}}(\mathbf{\theta})\big{)}^{2}$ (5)
@@ 3.1 Data set
Taking $m=250000$ to be the total number of data samples $\mathbf{g}=(x^{\mathrm{GT}},h^{\mathrm{GT}})^{T}$ used, a large population in the GT distribution is sampled from $G_{L}$ with $0.85m$ samples $\mathbf{g}_{0},\dots,\mathbf{g}_{0.85m-1}$ and a small population is sampled from $G_{s}$ with $0.15m$ samples $\mathbf{g}_{0.85m},\dots,\mathbf{g}_{m}$, without loss of generality.
@@ 4 Application to collider physics
The refinement protocol using the 2-stage training scheme introduced above is applied to a realistic example tailored to high-energy physics data collected at the LHC. The data set is described, choices regarding the training process are discussed, and the results of the refinement application follow.
@@ 4.1 Data set
Jets are notoriously difficult to model because their production mechanisms are only partly calculable with perturbative methods and because the energetic response of the detectors is a complicated function of the detector geometry, event activity, and radiation exposure. Considerable effort has been made to accurately model jets in state-of-the-art fullsim models such as Geant4-based [27, 28, 29] applications of ATLAS and CMS. Fullsim processing time per event ranges from 10s of second to a few minute per event, depending on the underlying physics process. In fastsim programs, such as ATLAS AtlFast3 [3], CMS FastSim [4, 5], the processing time is only a few seconds, but due to simplified assumptions in the calorimetry and tracking detectors, mismodeling of jets is often compounded. In parametric simulators, such as the Delphes framework [6], the processing time per event ranges in the milliseconds, but jet mismodeling arises due to simplified energy response parametrizations.
@@
We consider a set of jet substructure observables called N-subjettiness $\tau_{N}$ [30]. Ratios such as $\tau_{2}/\tau_{1}$ provide information on the compositeness of the jets to infer the unobservable particles from which they originated. A GT data set is generated based on proton-proton collisions at $\sqrt{s}=13\,\mathrm{TeV}$ using the Pythia 8.1 software package [31], consisting of events with a pair of top quarks in the final state. These events are then processed twice in parallel using Delphes, once with the default CMS detector implementation and treated as the fullsim data set, and once with a “flawed" implementation yielding the data set we treat as the fastsim. The flaw is introduced by setting the $\beta$ parameter [30] to 0.9 rather than 1.0, inducing an angle-dependent bias in the N-subjettiness, and setting the energy resolution of the electromagnetic calorimeter to a constant of 1% rather than the default values, which depend on transverse momentum $p_{T}$ and pseudorapidity $\eta$ [32, 33].
@@ 4.2 Training
$\mathbf{x^{\prime}}$ $\intercal\displaystyle=\begin{pmatrix}\frac{\tau_{2}}{\tau_{1}},&\frac{\tau_{3}}{\tau_{2}},&\frac{\tau_{4}}{\tau_{3}}\end{pmatrix}^{\intercal}$ (8)
$\mathbf{g}$ $\intercal\displaystyle=\begin{pmatrix}\frac{\tau_{2}^{GT}}{\tau_{1}^{GT}},&\frac{\tau_{3}^{GT}}{\tau_{2}^{GT}},&\frac{\tau_{4}^{GT}}{\tau_{3}^{GT}}\end{pmatrix}^{\intercal}\,.$ (9)
@@ 5 Conclusions
We have introduced a regression-based procedure for refining the final analysis variables produced by a fast simulation application (fastsim) to render them highly accurate with respect to cost-intensive full simulation (fullsim). The refinement model, which is based on a residual neural network, is trained by minimizing an ensemble loss (MMD), which evaluates the similarity between two unbinned multi-dimensional distributions. As an additional option, a secondary pair-based loss (MSE) is used in a 2-stage training prescription: one stage that combines the MMD and MSE losses, and a second stage which uses only the MMD.
Two examples have been explored for refinement, each considering the accuracy of features that are available to the network as well as hidden features which are not. An abstract example refines fastsim and fullsim proxy data sets defined analytically and synchronized at the level of the ground truth. We find that the single-stage MMD-only procedure performs well when evaluating the accuracy of the available feature; however, it falls short in describing the multidimensional target space which includes the hidden dimension. Introducing a 2-stage prescription that makes use of the MSE loss improves the accuracy, improving the description in the multidimensional domain. A second example based on a realistic model of collider physics at the CERN LHC examines the performance of the two prescriptions with a larger number of available and hidden features. Here, the 2-stage prescription leads to similar quality of refinement as the single-stage prescription.
Fast Perfekt makes comprehensive use of existing domain knowledge present in fastsim programs, minimizing the need for the network to learn most of the salient features and correlations present in the target data. The refinement acts on final variables or summary statistics of the simulation (i.e., jet properties) rather than intermediate representations (i.e., calorimeter shower hits). The utility of the MMD as a loss function for regression has been demonstrated, and the MMD is deemed suitable for refinement problems that target real-world rather than simulated data. In cases where the input and target samples can be synchronized at the level of the ground truth, the 2-stage Fast Perfekt prescription is preferred, as it constrains the network to a highly unique configuration by making use of the MSE loss. This increases the stability and overall accuracy of the refined simulation, and is considered ideal for tuning fastsim to match fullsim.